OpenStudy (anonymous):

Create a quadratic function, f(x), in vertex form. The a should be between 4 and –4, the h will be your birth month, and the k will be your birth day. Write your equation below.

OpenStudy (phi):

vertex form: \[ f(x) = a(x - h)^2 + k \]

OpenStudy (phi):

replace the letters a, h and k with numbers

OpenStudy (anonymous):

f(x) = 2(x-11)2 + 27 if my birthday is 11/27/96?

OpenStudy (anonymous):

@phi

OpenStudy (phi):

yes, but I would use ^ to show the 2 is an exponent f(x) = 2(x-11)^2 + 27

OpenStudy (anonymous):

Sweet, can you help me with a couple more based on this one?

OpenStudy (phi):

please make a new post

OpenStudy (anonymous):

Will do :)